OpenStudy (anonymous):

how do i get y by itself -3x-y=-10 4x-4y=8

OpenStudy (chris):

If you multiply both sides of an equation by the same number, you get an equivalent form of that equation. So in these cases, you want to multiply the top equation by one number, and the bottom equation by a different number, so that if you add the equations the X's cancel out. It's a lot like finding a lowest common denominator.

OpenStudy (chris):

In this case, if you multiple the top equation by 4, you get: 4(-3x-y) = 4(-10) - you have to multiply both sides of the equation by the same thing remember. The above results in: -12x -4y = -40

OpenStudy (chris):

so the next part is multiplying the bottom equation by "some number", such that the number of x's will cancel out with -12x

OpenStudy (chris):

in this case you can multiply the bottom equation by 3

OpenStudy (chris):

3*(4x-4y) = 3*8 12x - 12y = 24

OpenStudy (chris):

-12x -4y = -40 12x - 12y = 24 -16y = -16 y = 1

hero (hero):

Or you can simply add y to both sides while adding 10 to both sides in the first equation: So -3x-y=-10 becomes 3x + 10 = y

OpenStudy (chris):

Hehe true, I suppose I solved it

hero (hero):

I was just answering the question "how to get y by itself"

OpenStudy (chris):

yep understand.

OpenStudy (anonymous):

thankyou both :-)!